Diagonzalize the matrix A.if possible. That is, find an invertible matrix P and 1 3 3 Diagonalize the matrix A= - 3 - 5 -3 3 3 a diagonal matrix D such
Answer #1

Given matrix: ri ww -5 -3 3 1 The characteristic polynomial: 3 -3 3 -5- 3 1-1 2 + 4 -3-322 = -(4-1 69-74*+4) = -(4-1) (0+2) (For 22 = -2 A dz 1 -2 -3 I det av = xv So we have ous system of >> (4-73) 050 homogeneon equations we solve it by Gaussian El